# PDV: Prompt Directional Vectors for Zero-shot Composed Image Retrieval

### v2

@@ A.1 PDV Algorithm and Code
The PDV algorithm is given in Algorithm 1, and the code is shown in Figure 5. The implementation of PDV is very intuitive, and it could be easily integrated with any ZS-CIR approaches.
@@
Figure 5: Python function for calculating PDV features.
A.2 Additional Results
A.2.1 Additional Results
We also provide additional PDV-I results achieved on the validation set of the FashionIQ dataset, as shown in Table 4. PDV-I also achieved significant improvements over existing approaches that directly leverage image embeddings for retrieval.
Lastly, we provide a detailed visualization of the impact of $\alpha$/$\beta$ scaling on top-5 retrieval results. Figure 6 illustrates the performance of CIReVL with the ViT-B-32 CLIP model across three different datasets.
@@ A.2.1 Additional Results
Backbone Method $\alpha$ R@10 R@50 R@10 R@50 R@10 R@50 R@10 R@50
@@
SEARLE + PDV-I 24.48 42.30 18.79 38.47 25.91 47.02 23.03 42.60
CIReVL + PDV-I 28.95 45.88 29.00 49.13 34.22 56.09 30.72 50.37
Table 4: PDV-I performance on FashionIQ val datasets. † denotes that numbers are taken from the original paper.
A.2.2 Ablation Analysis
While Figure 3 in the main paper illustrates the effects of scaling factor $\alpha$ and fusion factor $\beta$ on Recall@5 performance across various PDV applications, Figures 7, 8, and 9 present complementary results for Recall@10 and Recall@50 metrics.
Figure 6: Visualisation of the impact of $\alpha$/$\beta$ scaling on top-5 retrieval results. CIReVL with ViT-B-32 Clip model is the baseline method used. Representative examples with prompts from three datasets: FashionIQ (left), CIRR (middle), and CIRCO (right) are shown at the top. Green and blue bounding boxes indicate true positives and near-true positives, respectively.
Figure 7: PDV-T: Impact of $\alpha$ scaling on Recall@10 (left) and Recall@50 (right) performance. Results shown for three baseline methods: CIReVL (top), Pic2Word (middle) and SEARLE (bottom).
Figure 8: PDV-I: Impact of $\alpha$ scaling on Recall@10 (left) and Recall@50 (right) performance. Results shown for three baseline methods: CIReVL (top), Pic2Word (middle) and SEARLE (bottom).
Figure 9: PDV-F: Impact of $\beta$ scaling on Recall@10 (left) and Recall@50 (right) performance. Results shown for three baseline methods: CIReVL (top), Pic2Word (middle) and SEARLE (bottom).

### v3

@@ A.4.2 PDV-I Results
The PDV algorithm is given in Algorithm 1, and the code is shown in Figure S10. The implementation of PDV is very intuitive, and it could be easily integrated with any ZS-CIR approaches.
@@
Figure S10: Python function for calculating PDV features.
Figure S11: Visualisation of the impact of $\alpha$$\beta$ scaling on top-5 retrieval results. CIReVL with ViT-B-32 Clip model is the baseline method used. Representative examples with prompts from three datasets: FashionIQ (left), CIRR (middle), and CIRCO (right) are shown at the top. Green and blue bounding boxes indicate true positives and near-true positives, respectively.
Figure S12: PDV-T: Impact of $\alpha$ scaling on Recall@10 (left) and Recall@50 (right) performance. Results shown for three baseline methods: CIReVL (top), Pic2Word (middle) and SEARLE (bottom).
Figure S13: PDV-I: Impact of $\alpha$ scaling on Recall@10 (left) and Recall@50 (right) performance. Results shown for three baseline methods: CIReVL (top), Pic2Word (middle) and SEARLE (bottom).
Figure S14: PDV-F: Impact of $\beta$ scaling on Recall@10 (left) and Recall@50 (right) performance. Results shown for three baseline methods: CIReVL (top), Pic2Word (middle) and SEARLE (bottom).
@@
Backbone Method $\alpha_{I}$ R@10 R@50 R@10 R@50 R@10 R@50 R@10 R@50
@@
SEARLE + PDV-I 2 18.25 31.84 18.49 39.17 21.32 37.74 19.35 36.25
CIReVL + PDV-I 2 28.95 45.88 29.00 49.13 34.22 56.09 30.72 50.37
Table S7: PDV-I performance on FashionIQ val datasets. † denotes that numbers are taken from the original paper.
Dataset CIRCO CIRR
Metric mAP@k Recall@k $R_{s}$@k
Arch Method $\alpha_{I}$ k=5 k=10 k=25 k=50 k=1 k=5 k=10 k=50 k=1 k=2 k=3
Table S8: PDV-I performance on CIRCO and CIRR test datasets. Note that the image-only approach utilizes the visual embedding of the reference image, whereas the text-only approach employs the text embedding of the prompt. Bold = best results, underline = second best. †Numbers from original paper.